# ETA: Evaluating Then Aligning Safety of Vision Language Models at Inference Time

### v1

@@ Abstract
Vision Language Models (VLMs) have become essential backbones for multimodal intelligence, yet significant safety challenges limit their real-world application. While textual inputs are often effectively safeguarded, adversarial visual inputs can easily bypass VLM defense mechanisms. Existing defense methods are either resource-intensive, requiring substantial data and compute, or fail to simultaneously ensure safety and usefulness in responses. To address these limitations, we propose a novel two-phase inference-time alignment framework, Evaluating Then Aligning (ETA): i) Evaluating input visual contents and output responses to establish a robust safety awareness in multimodal settings, and ii) Aligning unsafe behaviors at both shallow and deep levels by conditioning the VLMs’ generative distribution with an interference prefix and performing sentence-level best-of-$N$ to search the most harmless and helpful generation paths. Extensive experiments show that ETA outperforms baseline methods in terms of harmlessness, helpfulness, and efficiency, reducing the unsafe rate by 87.5% in cross-modality attacks and achieving 96.6% win-ties in GPT-4 helpfulness evaluation. The code is publicly available at https://github.com/DripNowhy/ETA.
@@ 1 Introduction
Vision Language Models (VLMs) (Achiam et al., 2023; Chen et al., 2023a; b; Bai et al., 2023; Liu et al., 2024a; b; Zhang et al., 2024b) have emerged as crucial multimodal intelligence backbones, offering unprecedented capabilities in processing and understanding both visual and textual information. These models are developed by integrating visual model backbones into pre-trained Large Language Models (LLMs), followed by visual instruction tuning. While VLMs have demonstrated excellent performance across various vision-language tasks, their real-world applications are significantly hindered by safety challenges. Zong et al. (2024) suggests that text-image pairs introduced during visual instruction tuning may contain unsafe content, potentially causing the model to forget safety mechanisms previously learned by the LLM backbone. Furthermore, research by Gong et al. (2023), Liu et al. (2023a), and Gou et al. (2024) indicates that the visual modality can easily bypass existing safety mechanisms, leading to harmful responses.
To ensure both harmlessness and helpfulness in VLM responses, current approaches can be categorized into fine-tuning-based and inference-based defenses (Jin et al., 2024). Fine-tuning-based methods, such as supervised fine-tuning (SFT) (Zong et al., 2024) and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Sun et al., 2023), aim to align models with human preferences but are resource-intensive, requiring extensive data and labor, and may compromise the model’s general capabilities (Zhang et al., 2024c; Dubey et al., 2024). Inference-based defenses attempt to leverage VLMs’ intrinsic safety mechanisms by using evaluators to judge response quality and ensure safe outputs based on predefined criteria (Wang et al., 2024b; Gou et al., 2024). However, existing methods often focus on evaluating textual inputs and outputs, while largely overlooking the crucial safety checks for visual inputs. This oversight can result in mis-evaluations in multimodal scenarios, allowing harmful content to slip through and ultimately weakening the model’s overall capabilities (Liu et al., 2024c; Jin et al., 2024). Furthermore, when unsafe behaviors are detected, current methods often rely on generic refusal responses, failing to offer helpful suggestions or relevant content based on the input (Pi et al., 2024). As a result, the responses become unhelpful and disconnected from the input topic.
@@ 2 Related Works
To enable VLMs to generate responses aligned with human preferences (e.g. harmless and helpful), approaches like reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Sun et al., 2023; Zhang et al., 2024c) or supervised fine-tuning on specialized datasets (Chen et al., 2024; Zong et al., 2024; Li et al., 2024c) are often employed. Other approaches aim to improve safety mechanisms by redesigning network architectures. For example, Bethany et al. (2024) and Liu et al. (2024c) introduce additional classifiers during training to assess the safety level and type of unsafe content, enhancing interpretability of model generation. Another line of work incorporates fine-tuned defense LLMs during inference to guide or correct model outputs, ensuring safer responses (Pi et al., 2024). However, these methods are resource-intensive, and the balance of harmful and helpful data in training sets can affect the model’s core capabilities. Furthermore, their safety capabilities are often limited to the specific domains represented in red-teaming data, resulting in weak generalization to other domains and adversarial attacks (Shayegani et al., 2023; Gou et al., 2024; Jin et al., 2024). In contrast, our approach requires no additional data or fine-tuning of model parameters, providing a plug-and-play solution to align VLMs.
@@
Inference-time alignment modifies the decoding strategy of language models to align output distributions with human preferences (Brown et al., 2024a; Zhang et al., 2024a). In LLM alignment, Khanov et al. (2024) and Li et al. (2024a) utilize reward models to score outputs and select the response with higher reward score based on predefined criteria. These methods avoid the instability associated with PPO training in RLHF (Andrychowicz et al., 2021; Zheng et al., 2023), while ensuring alignment with human preferences by sampling high-reward responses. Other techniques utilize self-evaluation of LLMs, employing the concept of LLM-as-a-Judge to evaluate candidate responses without introducing additional models (Xie et al., 2024; Brown et al., 2024b).
@@ 3 Preliminaries
To enable LLMs to understand visual information, mainstream methods generally involve two key steps: first, training a vision-language connector module, and second, fine-tuning the LLM’s backbone with various SFT image-text datasets (Dai et al., 2023; Liu et al., 2024a; b). During inference, for a given image-text pair $\{x_{I},x_{T}\}$, the vision encoder $\mathcal{C}$ first converts the image $x_{I}$ into a visual embedding $e_{I}$. The connector module $\mathcal{M}$ then projects $e_{I}$ into a continuous sub-space of the textual embedding space, which can serve as input to the LLMs. Similar to LLMs, VLMs generate responses by predicting the next token’s distribution in an autoregressive manner, continuing to generate tokens until a complete response is produced:
@@ 4 Methodology
Figure 2: Continuous visual token embeddings can bypass existing safety mechanisms that are primarily aligned with discrete textual token embeddings. To verify this hypothesis, we implemented a mapping that transforms continuous visual embeddings to their nearest discrete textual embeddings based on cosine similarity. This mapping results in a significant 6.79% reduction in the unsafe rate (USR) when evaluated on the SPA-VL Harm test set (Zhang et al., 2024c). Fig. 7 illustrates examples of these mapped textual tokens, demonstrating how offensive images are transformed into harmful tokens that can then be effectively addressed by the original safety mechanisms of LLM backbones.
@@ 4.1 Multimodal Evaluation
LLM backbones are typically aligned on discrete textual embeddings $\mathcal{E}_{\text{textual}}\subset\mathbb{R}^{d}$ (Devlin, 2018; Dubey et al., 2024). In contrast, the continuous visual embeddings $\mathcal{E}_{\text{visual}}\subset\mathbb{R}^{d}$ often appear away from all textual embeddings (Gong et al., 2023). As shown in Fig. 2, we implemented an alternating mapping, where continuous visual embeddings are mapped to their nearest textual embeddings (green guard flow). This method resulted in a significant 7% reduction in the unsafe rate (USR) compared to the standard VLM baseline (red bypass flow). We also show examples of mapped textual tokens in Fig. 7 and 8, where offensive images are mapped to related harmful words. These results provide direct evidence supporting our hypothesis that the bypassing of safety mechanisms in VLMs is primarily due to outliers in the embedding space, specifically those originating from the visual modality. Additionally, as evidenced by the red curve in Fig. 3(a), previous safety mechanisms built on LLMs fail in multimodal inputs. These insights inspired us to establish a new safety awareness for VLMs in multimodal settings, designed to safeguard both visual and textual information.
@@ 4.1.1 Pre-generation evaluator
Considering that many VLMs use CLIP-based vision towers (Chen et al., 2023b; Liu et al., 2024a; Zhang et al., 2024b), such as CLIP-ViT-L-336px [^1] https://huggingface.co/openai/clip-vit-large-patch14-336, this has inspired us to leverage the CLIP-score to develop semantic-level evaluation for visual inputs. Unlike the modality alignment in VLMs, CLIP models learn and align vision and text embeddings by maximizing the semantic similarity across modalities during training (Radford et al., 2021). Despite the continuous nature of visual embeddings, the safety of images can be determined by assessing the presence of unsafe semantic content.
@@ 4.1.2 Post-generation evaluator
Reward models (RMs) are trained on preference text datasets to evaluate the utility of responses and their alignment with human preferences (Li et al., 2024a). To compensate for the lack of visual modality in RM evaluation, we introduce a safety-specific input format (Appendix A.2), which compels the RM to evaluate responses based on both utility and safety criteria. In Fig. 4, we present the distribution of reward scores across different input formats. The results show that the safety-specific input format creates a more distinct separation between harmful and harmless responses compared to the vanilla format, allowing for more effective discrimination between safe and unsafe outputs. We define the reward score derived from our proposed safety-specific input format as the post-generation evaluation score, $\mathcal{S}_{\text{post}}$, calculated as:
@@ 4.2.2 Sentence-level best-of-$N$ seaching as deep alignment
where $\alpha$ balances safety ($\mathcal{S}_{\text{post}}(\cdot)$ in Eq. 4) and utility ($\mathcal{S}_{u}(\cdot,\cdot)$ in Eq. 5), and $O_{\leq i}$ represents the complete response up to the $i$-th sentence. For the first sentence with the interference prefix, $\alpha=0$ to ensure safety. For subsequent sentences, $\alpha=1/i$, as discussed in Section 4.2.1, to address the risk of harmful responses after the initial interference prefix.
@@ 5.1 Setups
We employed LLaVA-1.5-7B and 13B (Liu et al., 2024a), InternVL-Chat-1.0-7B (Chen et al., 2023b), and InternLM-XComposer-2.5-7B (Zhang et al., 2024b) as the VLM backbones. The textual RM used in ETA was ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a), which exhibits strong safety ability. For our ETA method, during the evaluation phase, we empirically set the thresholds to $\tau_{\text{pre}}=0.16$ in Eq. 3 and $\tau_{\text{post}}=0.06$ in Eq. 4. In the alignment phase, we generated $N=5$ candidate responses per sentence. All experiments were conducted on an NVIDIA RTX A6000 platform. The prompts used during pre- and post-generation evaluations are detailed in Appendix A.1 and A.2.
@@
We focus on two main categories of benchmarks to evaluate VLM capabilities: safety and helpfulness. For safety, we assess ETA using multimodal safety datasets, including SPA-VL Harm (Zhang et al., 2024c), MM-SafetyBench (Liu et al., 2023a), FigStep (Gong et al., 2023), Unconstrained attack (Qi et al., 2024a), and the text attack benchmark AdvBench (Zou et al., 2023). Following the methodology of Zhang et al. (2024c), we use the LLM safety guard-based Unsafe Rate (USR) as the primary evaluation metric, which measures the proportion of unsafe responses generated by the model. Additionally, in line with Zong et al. (2024) and Wang et al. (2024b), we compute the target-string-based Attack Success Rate (ASR) as a complementary metric. For helpfulness, we selected several common comprehensive benchmarks and VQA datasets, such as $\text{SQA}^{I}$ (ScienceQA-IMG) (Lu et al., 2022), VQAv2 (Goyal et al., 2017b), TextVQA (Singh et al., 2019), MME (Fu et al., 2023), and MMBench (Liu et al., 2023b), to evaluate the general capabilities of VLMs. Additionally, we used GPT-4-Turbo to assess the helpfulness of model outputs on the SPA-VL Help dataset (Zhang et al., 2024c). Further details on evaluation metrics and benchmarks are provided in Appendix A.3 and A.4.
@@
Given that ETA requires no additional data or fine-tuning, we primarily compare it against existing inference-time method, ECSO (Gou et al., 2024). Additionally, to demonstrate that our approach can reduce VLM safety issues while maintaining output usefulness, we also compare it with fine-tuned methods like Posthoc-LoRA and Mixed-LoRA on VLGuard in the helpfulness evaluation (Zong et al., 2024).
@@ 5.2 Results
Table 1: USR evaluation across multiple safety benchmarks. Our method significantly reduces unsafe responses to malicious inputs across four different VLM backbones. Under suffix adversarial attacks and cross-modality attack, ETA demonstrates superior performance, while ECSO fails to generate safe responses under these conditions.
@@
Compared to ECSO, ETA significantly reduces USR across all benchmarks. Notably, when facing adversarial attacks, the minimal difference in USR between ECSO and the VLM backbone indicates that ECSO does not truly safeguard the model but rather limits the impact of the input image on the model’s safety. In contrast, our method remains unaffected by these attacks, reducing LLaVA-1.5-13B’s USR on cross-modality attack by 87.5%, compared to ECSO’s mere 5% reduction. Detailed results including that ETA achieves outstanding safety performance on text-only benchmarks are reported in Table 7, as well as its effectiveness on target-string-based metrics in Table 6.
@@
As shown in Table 2, compared to fine-tuned methods, inference-time approaches have a smaller impact on the VLM backbones in the Comprehensive and VQA benchmarks. Furthermore, our ETA does not diminish the backbone’s capabilities in any of the five benchmarks. The only exception is TextVQA, where ETA reduces accuracy by just 0.05%, still better than other baselines. This indicates that ETA provides more reliable assessments during the evaluating phase, ensuring that the model’s general abilities remain unaffected. Results from evaluating additional VLM backbones are provided in Table 8 of Appendix B.4.
@@
Inference efficiency is vital for inference-time alignment methods. Given the common usage scenarios of VLMs, we compared the inference time of ETA and ECSO on the comprehensive MMB and $\text{SQA}^{I}$ benchmark, with the results provided in Table 4. It can be observed that ETA increases the inference time per generation by only 0.1 seconds compared to the VLM backbone, whereas ECSO adds an average of 0.39 seconds, almost 4 times the increase of ETA. This is because ECSO’s self-evaluation struggles to accurately assess response safety, leading to an excessive number of unnecessary alignment steps during generation. In contrast, ETA provides accurate evaluation, preserving VLMs’ general ability and avoiding the extra generation overhead.

### v2

@@ Abstract
Vision Language Models (VLMs) have become essential backbones for multimodal intelligence, yet significant safety challenges limit their real-world application. While textual inputs can often be effectively safeguarded, adversarial visual inputs can easily bypass VLM defense mechanisms. Existing defense methods are either resource-intensive, requiring substantial data and compute, or fail to simultaneously ensure safety and usefulness in responses. To address these limitations, we propose a novel two-phase inference-time alignment framework, Evaluating Then Aligning (ETA): i) Evaluating input visual contents and output responses to establish a robust safety awareness in multimodal settings, and ii) Aligning unsafe behaviors at both shallow and deep levels by conditioning the VLMs’ generative distribution with an interference prefix and performing sentence-level best-of-$N$ to search the most harmless and helpful generation paths. Extensive experiments show that ETA outperforms baseline methods in terms of harmlessness, helpfulness, and efficiency, reducing the unsafe rate by 87.5% in cross-modality attacks and achieving 96.6% win-ties in GPT-4 helpfulness evaluation. The code is publicly available at https://github.com/DripNowhy/ETA.
@@ 1 Introduction
Vision Language Models (VLMs) (Achiam et al., 2023; Chen et al., 2023a; b; Bai et al., 2023; Liu et al., 2024a; c; Zhang et al., 2024b) have emerged as crucial multimodal intelligence backbones, offering unprecedented capabilities in processing and understanding both visual and textual information. These models are developed by integrating visual model backbones into pre-trained Large Language Models (LLMs), followed by visual instruction tuning. While VLMs have demonstrated excellent performance across various vision-language tasks, their real-world applications are significantly hindered by safety challenges (Tu et al., 2023). Zong et al. (2024) suggests that text-image pairs introduced during visual instruction tuning may contain unsafe content, potentially causing the model to forget safety mechanisms previously learned by the LLM backbone. Furthermore, research by Gong et al. (2023), Liu et al. (2023a), and Gou et al. (2024) indicates that the visual modality can easily bypass existing safety mechanisms, leading to harmful responses.
To ensure both harmlessness and helpfulness in VLM responses, current approaches can be categorized into fine-tuning-based and inference-based defenses (Jin et al., 2024). Fine-tuning-based methods, such as supervised fine-tuning (SFT) (Zong et al., 2024) and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Sun et al., 2023), aim to align models with human preferences but are resource-intensive, requiring extensive data and labor, and may compromise the model’s general capabilities (Zhang et al., 2024c; Dubey et al., 2024). Inference-based defenses attempt to leverage VLMs’ intrinsic safety mechanisms by using evaluators to judge response quality and ensure safe outputs based on predefined criteria (Wang et al., 2024b; Gou et al., 2024). However, existing methods often focus on evaluating textual inputs and outputs, while largely overlooking the crucial safety checks for visual inputs. This oversight can result in mis-evaluations in multimodal scenarios, allowing harmful content to slip through and ultimately weakening the model’s overall capabilities (Liu et al., 2024d; Jin et al., 2024). Furthermore, when unsafe behaviors are detected, current methods often rely on generic refusal responses, failing to offer helpful suggestions or relevant content based on the input (Pi et al., 2024). As a result, the responses become unhelpful and disconnected from the input topic.
@@ 2 Related Works
To enable VLMs to generate responses aligned with human preferences (e.g. harmless and helpful), approaches like reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Sun et al., 2023; Zhang et al., 2024c) or supervised fine-tuning on specialized datasets (Chen et al., 2024; Zong et al., 2024; Li et al., 2024d) are often employed. Other approaches aim to improve safety mechanisms by redesigning network architectures. For example, Bethany et al. (2024) and Liu et al. (2024d) introduce additional classifiers during training to assess the safety level and type of unsafe content, enhancing interpretability of model generation. Another line of work incorporates fine-tuned defense LLMs during inference to guide or correct model outputs, ensuring safer responses (Pi et al., 2024). However, these methods are resource-intensive, and the balance of harmful and helpful data in training sets can affect the model’s core capabilities. Furthermore, their safety capabilities are often limited to the specific domains represented in red-teaming data, resulting in weak generalization to other domains and adversarial attacks (Shayegani et al., 2023; Gou et al., 2024; Jin et al., 2024). In contrast, our approach requires no additional data or fine-tuning of model parameters, providing a plug-and-play solution to align VLMs.
@@
Inference-time alignment modifies the decoding strategy of language models to align output distributions with human preferences (Brown et al., 2024a; Zhang et al., 2024a). In LLM alignment, Khanov et al. (2024) and Li et al. (2024b) utilize reward models to score outputs and select the response with higher reward score based on predefined criteria. These methods avoid the instability associated with PPO training in RLHF (Andrychowicz et al., 2021; Zheng et al., 2023), while ensuring alignment with human preferences by sampling high-reward responses. Other techniques utilize self-evaluation of LLMs, employing the concept of LLM-as-a-Judge to evaluate candidate responses without introducing additional models (Xie et al., 2024; Brown et al., 2024b).
@@ 3 Preliminaries
To enable LLMs to understand visual information, mainstream methods generally involve two key steps: first, training a vision-language connector module, and second, fine-tuning the LLM’s backbone with various SFT image-text datasets (Dai et al., 2023; Liu et al., 2024a; c). During inference, for a given image-text pair $\{x_{I},x_{T}\}$, the vision encoder $\mathcal{C}$ first converts the image $x_{I}$ into a visual embedding $e_{I}$. The connector module $\mathcal{M}$ then projects $e_{I}$ into a continuous sub-space of the textual embedding space, which can serve as input to the LLMs. Similar to LLMs, VLMs generate responses by predicting the next token’s distribution in an autoregressive manner, continuing to generate tokens until a complete response is produced:
@@ 4 Methodology
Figure 2: Continuous visual token embeddings can bypass existing safety mechanisms that are primarily aligned with discrete textual token embeddings. To verify this hypothesis, we implemented a mapping that transforms continuous visual embeddings to their nearest discrete textual embeddings based on cosine similarity. This mapping results in a significant 7% reduction in the unsafe rate (USR) when evaluated on the SPA-VL Harm test set (Zhang et al., 2024c) (We report more results on four VLM baselines and two datasets in Appendix B.1). Fig. 8 illustrates examples of these mapped textual tokens, demonstrating how offensive images are transformed into harmful tokens that can then be effectively addressed by the original safety mechanisms of LLM backbones.
@@ 4.1 Multimodal Evaluation
LLM backbones are typically aligned on discrete textual embeddings $\mathcal{E}_{\text{textual}}\subset\mathbb{R}^{d}$ (Devlin, 2018; Dubey et al., 2024). In contrast, the continuous visual embeddings $\mathcal{E}_{\text{visual}}\subset\mathbb{R}^{d}$ often appear away from all textual embeddings (Gong et al., 2023). As shown in Fig. 2, we implemented an alternating mapping, where continuous visual embeddings are mapped to their nearest textual embeddings (green guard flow). This method resulted in a significant 7% reduction in the unsafe rate (USR) compared to the standard VLM baseline (red bypass flow). We also show examples of mapped textual tokens in Fig. 8 and 9, where offensive images are mapped to related harmful words. These results provide direct evidence supporting our hypothesis that the bypassing of safety mechanisms in VLMs is primarily due to outliers in the embedding space, specifically those originating from the visual modality. Additionally, as evidenced by the red curve in Fig. 3(a), previous safety mechanisms built on LLMs fail in multimodal inputs. These insights inspired us to establish a new safety awareness for VLMs in multimodal settings, designed to safeguard both visual and textual information.
@@ 4.1.1 Pre-generation evaluator
Considering that many VLMs use CLIP-based vision towers (Chen et al., 2023b; Liu et al., 2024a; Zhang et al., 2024b), such as CLIP-ViT-L-336px [^1] https://huggingface.co/openai/clip-vit-large-patch14-336, this highlights that continuous visual embeddings are essentially derived from the pre-trained CLIP vision encoder. This insight inspired us to leverage the CLIP score to propose a semantic-level evaluation method for visual inputs. Unlike the modality alignment in VLMs, CLIP models learn and align vision and text embeddings by maximizing the semantic similarity across modalities during training (Radford et al., 2021). Despite the continuous nature of visual embeddings, the safety of images can be determined by assessing the presence of unsafe semantic content.
@@ 4.1.2 Post-generation evaluator
Reward models (RMs) are trained on preference text datasets to evaluate the utility of responses and their alignment with human preferences (Li et al., 2024b). To compensate for the lack of visual modality in RM evaluation, we introduce a safety-specific input format (Appendix A.2), which compels the RM to evaluate responses based on both utility and safety criteria. In Fig. 4, we present the distribution of reward scores across different input formats. The results show that the safety-specific input format creates a more distinct separation between harmful and harmless responses compared to the vanilla format, allowing for more effective discrimination between safe and unsafe outputs. We define the reward score derived from our proposed safety-specific input format as the post-generation evaluation score, $\mathcal{S}_{\text{post}}$, calculated as:
@@ 4.2.2 Sentence-level best-of-$N$ seaching as deep alignment
[…]
@@ 5.1 Setups
We employed LLaVA-1.5-7B and 13B (Liu et al., 2024a), LLaVA-NeXT-8B (Liu et al., 2024b), LLaVA-OneVision-7B-Chat (Li et al., 2024a), InternVL-Chat-1.0-7B (Chen et al., 2023b), InternLM-XComposer-2.5-7B (Zhang et al., 2024b), and Llama3.2-11B-Vision-Instruct (Dubey et al., 2024) as the VLM backbones. The textual RM used in ETA was ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a), which exhibits strong safety ability. For our ETA method, during the evaluation phase, we empirically set the thresholds to $\tau_{\text{pre}}=0.16$ in Eq. 3 and $\tau_{\text{post}}=0.06$ in Eq. 4. In the alignment phase, we generated $N=5$ candidate responses per sentence. All experiments were conducted on an NVIDIA RTX A6000 platform. The prompts used during pre- and post-generation evaluations are detailed in Appendix A.1 and A.2.
@@
We focus on two main categories of benchmarks to evaluate VLM capabilities: safety and helpfulness. For safety, we assess ETA using multimodal safety datasets, including SPA-VL Harm (Zhang et al., 2024c), MM-SafetyBench (Liu et al., 2023a), FigStep (Gong et al., 2023), Unconstrained attack (Qi et al., 2024a), and the text attack benchmark AdvBench (Zou et al., 2023). Following the methodology of Zhang et al. (2024c), we use the LLM safety guard-based Unsafe Rate (USR) as the primary evaluation metric, which measures the proportion of unsafe responses generated by the model. Additionally, in line with Zong et al. (2024) and Wang et al. (2024b), we compute the target-string-based Attack Success Rate (ASR) as a complementary metric. For helpfulness, we selected several common comprehensive benchmarks and VQA datasets, such as $\text{SQA}^{I}$ (ScienceQA-IMG) (Lu et al., 2022), VQAv2 (Goyal et al., 2017b), TextVQA (Singh et al., 2019), MME (Fu et al., 2023), MMBench (Liu et al., 2023b), and MMMU-Pro (Yue et al., 2024b) to evaluate the general capabilities of VLMs. Additionally, we used GPT-4-Turbo to assess the helpfulness of model outputs on the SPA-VL Help dataset (Zhang et al., 2024c). Further details on benchmarks and evaluation metrics are provided in Appendix A.3 and A.4.
@@
Given that ETA requires no additional data or fine-tuning, we primarily compare it against existing inference-time method, ECSO (Gou et al., 2024). Additionally, to demonstrate that our approach can reduce VLM safety issues while maintaining output usefulness, we also compare it with fine-tuned methods like Posthoc-LoRA and Mixed-LoRA on VLGuard (Zong et al., 2024) in the helpfulness evaluation.
@@ 5.2 Results
Table 1: USR evaluation across multiple safety benchmarks. Our method significantly reduces unsafe responses to malicious inputs across four different VLM backbones. Under suffix adversarial attacks and cross-modality attack, ETA demonstrates superior performance, while ECSO fails to generate safe responses under these conditions. Results on three more recent VLMs are provided in Table 7.
@@
Compared to ECSO, ETA significantly reduces USR across all benchmarks. Notably, when facing adversarial attacks, the minimal difference in USR between ECSO and the VLM backbone indicates that ECSO does not truly safeguard the model but rather limits the impact of the input image on the model’s safety. In contrast, our method remains unaffected by these attacks, reducing LLaVA-1.5-13B’s USR on cross-modality attack by 87.5%, compared to ECSO’s mere 5% reduction. Additionally, the results in Table 7 show that even for recent VLMs like Llama-3.2-Vision, which have undergone multimodal safety alignment, ETA further strengthens safeguarding, delivering exceptional safety performance on challenging cases. More results, including ETA’s strong performance on text-only benchmarks and its effectiveness on target-string-based metrics are reported in Table 9 and Table 8.
@@
As shown in Table 2, compared to fine-tuned methods, inference-time approaches have a smaller impact on the VLM backbones in the Comprehensive and VQA benchmarks. Furthermore, our ETA does not diminish the backbone’s capabilities in any of the five benchmarks. The only exception is TextVQA, where ETA reduces accuracy by just 0.05%, still better than other baselines. This indicates that ETA provides more reliable assessments during the evaluating phase, ensuring that the model’s general abilities remain unaffected. We present a performance comparison of various methods on LLaVA-1.5-13B in Table 10 and further evaluate our ETA on the challenging task MMMU-Pro, with results provided in Table 11 of Appendix B.6.
@@
Inference efficiency is vital for inference-time alignment methods. Given the common usage scenarios of VLMs, we compared the inference time of ETA and ECSO on the comprehensive MMB and $\text{SQA}^{I}$ benchmark, with the results provided in Table 4. It can be observed that ETA increases the inference time per generation by only 0.1 seconds compared to the VLM backbone, whereas ECSO adds an average of 0.39 seconds, almost 4 times the increase of ETA. This is because ECSO’s self-evaluation struggles to accurately assess response safety, leading to an excessive number of unnecessary alignment steps during generation. In contrast, ETA provides accurate evaluation, preserving VLMs’ general ability while avoiding the extra generation overhead.